OpenStudy (anonymous):

what are the factors of 125-y3

OpenStudy (lgbasallote):

difference of two cubes 125 = 5^3 y^3 = y^3 difference of two cubes is a^3 - b^3 = (a-b)(a^2 + ab + b^2)

OpenStudy (lgbasallote):

@benzonrambayon are you there?

OpenStudy (anonymous):

I don't think so..@lgbasallote maybe he/she fell asleep